Fermi National Laboratory
 

Operation Summary for June 22, 2011 - June 24, 2011

 

Wednesday June 22

The day shift began with Operations monitoring the Tevatron (TEV) store 8830 (210.58E30) and the Recycler stash (79.1E10), with the Antiproton Source (Pbar) stacking (18.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off, with the MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations established beam to NuMI at 9:23 AM; MI and Booster personnel tuned their machines to accommodate the NuMI beam. Operations lowered the beam intensity to MiniBooNE due to beam position problems. At 10:44 AM, the NuMI beam permit tripped off when a power supply (E:V118) tripped off; an EE Support person investigated and changed out a control board. NuMI resumed taking beam at 11:09 AM. At 11:16 AM, the NuMI horn tripped off twice; it reset both times. E:V118 continued to give NuMI problems and at 3:08 PM, with permission from NuMI, operators lowered a trip limit for the power supply. At 3:24 PM, the NuMI beam permit tripped off again, for the first of two times in two minutes, due to a quadrupole magnet power supply’s (E:Q107) output going low; it reset. Operators reported short downtimes due to problems with Linac and Booster RF stations (LRF1 and BRF7).

During the evening shift, at 5:05 PM, the NuMI beam permit tripped off due to E:V118; it reset. Operations disabled beam to T-978 at 6:06 PM. Operations terminated store 8830 at 7:04 PM. Operators prepared CDF and D0 for accesses. A LRF2 trip kept beam off for about 25 minutes. CDF and D0 came out of access and operators put the TEV through a dry squeeze at 10:23 PM. A TEV RF station (TRF4) tripped off on a LCW flow indication during the dry squeeze; operators called in a TEV RF person who reported at 11:22 PM that he had changed out a filter for the TRF4 LCW system.

Thursday June 23

The midnight shift began with Operations monitoring the stash (395.1E10), with Pbar stacking (59.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators put the TEV through a wet squeeze at 12:07 AM. Operations put the TEV into HEP shot setup at 12:32 AM and loaded protons and antiprotons at 1:17 AM, but the beam was aborted by D0 due to bad beam positions. Operators notified the TEV Coordinator about the trip at 1:52 AM. Operations established beam to T-978 at 4:37 AM. Operators put the TEV through at wet squeeze at 5:30 AM. Operations put the TEV into shot setup at 5:57 AM and established store 8835 at 7:24 AM with an average initial luminosity (AIL) of 202.2E30.

During the day shift, at 8:32 AM, operators reported that a couple of feed down sextupole magnet power supplies tripped off while they were tuning and caused the loss of about 2.5E30 of luminosity; operators were able to get the supplies back on. Operations sent one hour of beam to MTA (MuCool) starting at 8:52 AM. At 2 PM, operators reported problems with Booster beam positions. Booster personnel reported that all of the Booster cycles were affected by what ever was happening. Operation turned off MiniBooNE beam during the Booster beam investigations. Operators reported a short down time due to LRF3.

During the evening shift, at 4:53 PM, operators reported about 30 minutes of problems with LRF5. A RF person reported finding a loose connector on BRF5 at 5:46 PM. Operators brought MiniBooNE back online. Operations disabled beam to T-978 at 6:01 PM. Operations terminated store 8835 at 11:08 PM and put the TEV into shot setup.

Friday June 24

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (298.7E10), with Pbar stacking (13.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 8836 at 12:21 AM with an AIL of 391E30. At 1:26 AM, operators reported that BRF7 tripped off for the first of five times over 3.5 hours; it reset. Operations established beam to T-978 at 4:16 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow the following jobs today: Pbar will shutdown for about two hours so that FESS personnel can repair a vault exhaust fan; Booster will be down while RF personnel work on BRF5; MTA personnel will conduct a power supply test; and NuMI personnel will conduct a target scan this afternoon.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/24/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory